# Probabilistic Kernel Function for Fast Angle Testing

### v1

@@ Abstract
In this paper, we study the angle testing problem in high-dimensional Euclidean spaces and propose two projection-based probabilistic kernel functions, one designed for angle comparison and the other for angle thresholding. Unlike existing approaches that rely on random projection vectors drawn from Gaussian distributions, our approach leverages reference angles and employs a deterministic structure for the projection vectors. Notably, our kernel functions do not require asymptotic assumptions, such as the number of projection vectors tending to infinity, and can be both theoretically and experimentally shown to outperform Gaussian-distribution-based kernel functions. We further apply the proposed kernel function to Approximate Nearest Neighbor Search (ANNS) and demonstrate that our approach achieves a 2.5X $\sim$ 3X higher query-per-second (QPS) throughput compared to the state-of-the-art graph-based search algorithm HNSW.
@@ 1 Introduction
Vector-based similarity search is a core problem with broad applications in machine learning, data mining, and information retrieval. It involves retrieving data points in high-dimensional space that are most similar to a given query vector based on a specific similarity measure. This task is central to many downstream applications, including nearest neighbor classification, recommendation systems, clustering, anomaly detection, and large-scale information retrieval. However, the high dimensionality of modern datasets makes efficient similarity search particularly challenging, highlighting the need for fast and scalable vector computation techniques.
@@
Among the various similarity measures for high-dimensional vectors, the $\ell_{2}$ norm, cosine distance, and inner product are the most commonly used in practice. As discussed in Norm_range_LSH ; norm-explicit-quantization ; peos , it is often possible to pre-compute and store the norms of vectors in advance, allowing these measures to be reduced to the computation of the angle (inner product) between two normalized vectors, thereby highlighting the central role of angle computation. On the other hand, in many real-world scenarios, we are not concerned with the exact value of the angle itself but rather with the outcome of an angle-based comparison, which is referred to as the angle testing. Specifically, given a query vector $\bm{q}$ and data vectors $\bm{v_{1}}$, $\bm{v_{2}}$, $\bm{v}$ on sphere $\mathbb{S}^{d-1}$, typical operations include comparing $\left\langle\bm{q},\bm{v_{1}}\right\rangle$ and $\left\langle\bm{q},\bm{v_{2}}\right\rangle$, or determining whether $\left\langle\bm{q},\bm{v}\right\rangle$ exceeds a certain threshold. These operations, however, require computing exact inner products, which have a cost of $O(d)$ per comparison and become expensive in high dimensions. To address this, we aim to design a computation-efficient probabilistic kernel function $K$ that can approximate these comparisons with reduced cost and high success probability. In particular, we focus on the following two problems:
Problem 1.1
@@
(Probabilistic kernel function for comparison) Given $\bm{q}$, $\bm{v_{1}}$ and $\bm{v_{2}}$ on $\mathbb{S}^{d-1}$, where $\left\langle\bm{q},\bm{v_{1}}\right\rangle>\left\langle\bm{q},\bm{v_{2}}\right\rangle$, how can we design a probabilistic kernel function $K(\cdot,\cdot):\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}\rightarrow\text{RV}$, where RV denotes the set of real-valued random variables, such that (1) the computation of $K(\bm{q},\bm{v})$ does not rely on the computation of $\left\langle\bm{q},\bm{v}\right\rangle$, and (2) $\mathbb{P}[K(\bm{q},\bm{v_{1}})>K(\bm{q},\bm{v_{2}})]$ is as high as possible?
Problem 1.2
(Probabilistic kernel function for thresholding) Given an arbitrary pair of normalized vectors $(\bm{q},\bm{v})$ with an angle $\phi$, and an angle threshold $\theta$, how can we design a probabilistic kernel function $K(\cdot,\cdot):\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}\rightarrow\text{RV}$ such that (1) the computational complexity of $K(\bm{q},\bm{v})$ is significantly lower than that of computing the exact inner product $\left\langle\bm{q},\bm{v}\right\rangle$, and (2) $K(\bm{q},\bm{v})$ can reliably determine whether $\phi$ is smaller than $\theta$ with a high success probability?
@@
Both problems have various applications. The goal of Problem 1.1 aligns with that of CEOs ceos under cosine distance (we will postpone the general case of inner product to Sec. 4), allowing the designed kernel function to be applied to tasks where CEOs is effective, such as Maximum Inner Product Search (MIPS) ceos , filtering of NN candidates Falconn++ , DBSCAN sDBSCAN , and more. On the other hand, the goal of Problem 1.2 is similar to that of PEOs peos , making the corresponding kernel function well-suited for probabilistic routing tests in similarity graphs, which have demonstrated significant performance improvements over original graphs, such as HNSW hnsw . In Sec. 4, we will further elaborate on the applications of these two probabilistic kernel functions.
@@
Despite addressing different tasks, all the techniques ceos ; Falconn++ ; sDBSCAN ; peos mentioned above use Gaussian distribution to generate projection vectors and are built upon a common statistical result, as follows.
Lemma 1.3
@@
(Theorem 1 in ceos ) Given two vectors $\bm{v}$, $\bm{q}$ on $\mathbb{S}^{d-1}$, and $m$ random vectors $\{\bm{u}_{i}\}^{m}_{i=1}\sim\mathcal{N}(0,I^{d})$, and $m$ is sufficiently large, assuming that $\bm{u}_{\mathrm{max}}={\mathop{\rm argmax}_{\bm{u_{i}}}}|\bm{q}^{\top}\bm{u_{i}}|$, we have:
@@
However, Lemma 1.3 has a significant theoretical limitation: the relationship (1) relies on the assumption that the number of projection vectors $m$ tends to infinity. Since the evaluation time of projection vectors depends on $m$, $m$ cannot be very large in practice. Moreover, since Falconn++ ; sDBSCAN ; peos all used Lemma 1.3 to derive their respective theoretical results, these results are also affected by this limitation, and the impact of $m$ becomes even harder to predict in these applications.
The starting point of our research is to overcome this limitation, and we make the following two key observations. (1) The Gaussian distribution used in Lemma 1.3 is not essential. In fact, the only factor determining the estimation accuracy of $\bm{q}^{\top}\bm{v}$ is the reference angle, that is, the angle between $\bm{q}$ and $\bm{u_{\text{max}}}$. (2) By introducing a random rotation matrix, the reference angle becomes dependent on the structure of the projection vectors and is predictable.
Based on these two observations, we design new probabilistic kernel functions to solve problems 1.1 and 1.2. Specifically, the contributions of this paper are summarized as follows.
(1) The proposed kernel functions $K^{1}_{S}$ and $K^{2}_{S}$ (see eq. (2) and eq. (3)) rely on a reference-angle-based probabilistic relationship between angles in high-dimensional spaces and projected values. Compared with eq. (1), the new relationship (see relationship (4)) is deterministic without dependence on asymptotic condition. By theoretical analysis, we show that, the proposed kernel functions are effective for the Problems 1.1 and 1.2, and the smaller the reference angle is, the more accurate the kernel functions are (see Lemmas 3.2 and 3.3 ).
(2) To minimize the reference angle, we study the structure of the configuration of projection vectors (see Sec. 3.2). We propose two structures (see Alg. 1 and Alg. 2) that perform better than purely random projection (see Alg. 3), and we establish the relationship between the reference angle and the proposed structures (see Lemma 3.5 and Fig. 4).
(3) Based on $K^{1}_{S}$, we propose a random-projection technique KS1 which can be used for CEOs-based tasks ceos ; Falconn++ ; sDBSCAN . Based on $K^{2}_{S}$, we introduce a new routing test called the KS2 test which can be used to accelerate the graph-based Approximate Nearest Neighbor Search (ANNS) peos (see Sec. 4).
(4) We experimentally show that KS1 provides a slight accuracy improvement (up to 0.8%) over CEOs. In the ANNS task, we demonstrate that HNSW+KS2 improves the query-per-second (QPS) of the state-of-the-art approach HNSW+PEOs peos by 10% – 30%, along with a 5% reduction in index size.
@@ 2 Related Work
Due to space limitations, we focus on random projection techniques that are closely related to this work. An illustration comparing the proposed random projection technique with others can be found in Sec. B.1 and Fig. 3 in Appendix. Since the proposed kernel function is also used in similarity graphs for ANNS, a comprehensive discussion of ANNS solutions is provided in Appendix B.2.
In high-dimensional Euclidean spaces, the estimation of angles via random-projection techniques, especially Locality Sensitive Hashing (LSH) IndykM98:approximate-nearest-neighbor ; Andoni2004:e2lsh ; AndoniI08:near-optimal-hashing , has a relatively long history. One classical LSH-based technique is SimHash simhash , whose basic idea is to generate multiple hyperplanes and partition the original space into many cells such that two vectors falling into the same cell are likely to have a small angle between them. In falconn , the authors proposed a different LSH-based technique called Falconn for angular distance, whose basic idea is to find the closest or furthest projection vector to the data vector and record this projection vector as a hash value, leading to better search performance than SimHash. Later, the authors in ceos employed Concomitants of Extreme Order Statistics (CEOs) to identify the projection with the largest or smallest inner product with the data vector, as shown in Lemma 1.3, and recorded the corresponding maximum or minimum projected value to obtain a more accurate estimation than using a hash value alone Falconn++ . Notably, CEOs can be used not only for angular distance but also for inner product estimation.
Due to its ease of implementation, CEOs has been employed in several similarity search tasks ceos ; falconn ; sDBSCAN , as mentioned in Sec. 1. Additionally, CEOs has been used to accelerate similarity graphs, which are among the leading structures for Approximate Nearest Neighbor Search (ANNS). In peos , by swapping the roles of query and data vectors in CEOs, the authors introduced a space-partitioning technique and proposed the PEOs test, which can be used to compare the objective angle with a fixed threshold under probabilistic guarantees. This test was incorporated into the routing mechanisms of similarity graphs and achieved significant search performance improvements over original graph structures like HNSW hnsw and NSSG nssg .
@@ 3 Two Probabilistic Kernel Functions
3.1 Reference-Angle-Based Design
In Sec. 3.1, we aim to propose probabilistic kernel functions for Problems 1.1 and 1.2. First, we introduce some notation. Let $\mathbb{R}^{d}$ be the ambient vector space. Define $H\in SO(d)\in\mathbb{R}^{d\times d}$ as a random rotation matrix (Note that the definition here differs from that in falconn , where the so-called random rotation matrix is actually a matrix with i.i.d. Gaussian entries), and let $S=[\bm{u_{1}},\bm{u_{2}},\ldots,\bm{u_{m}}]\in\mathbb{R}^{d\times m}$ be an arbitrary fixed set of $m$ points on the unit sphere $\mathbb{S}^{d-1}$. For any vector $\bm{v}\in\mathbb{S}^{d-1}$, define the reference vector of $\bm{v}$ with respect to $S$ as $Z_{S}(\bm{v})=\mathop{\rm argmax}_{\bm{u}\in S}\left\langle\bm{u},\bm{v}\right\rangle$. Let $A_{S}(\bm{v})$ denote the cosine of the reference angle with respect to $Z_{S}(\bm{v})$, that is, $A_{S}(\bm{v})=\left\langle\bm{v},Z_{S}(\bm{v})\right\rangle$ Next, we introduce two probabilistic kernel functions $K^{1}_{S}(\cdot,\cdot)$ and $K^{2}_{S}(\cdot,\cdot)$ as follows, where $K^{1}_{S}(\cdot,\cdot)$ corresponds to Problem 1.1 and $K^{2}_{S}(\cdot,\cdot)$ corresponds to Problem 1.2.
@@ 3.1 Reference-Angle-Based Design
Remarks. (1) (Computational efficiency) Both functions are computationally efficient. For $K^{1}_{S}(\bm{q},\bm{v})$, $HS$ and $\left\langle\bm{v},H\bm{u_{i}}\right\rangle$ ($1\leq i\leq m$) can be pre-computed. Hence, for all $\bm{v}$’s, we only need to determine $Z_{HS}(\bm{q})$ online, which requires cost $O(dm)$ (This cost can be further reduced by a reasonable choice of $S$. See Sec. 3.2). For $K^{2}_{S}(\bm{q},\bm{v})$, $A_{S}(H\bm{v})$ and $Z_{S}(H\bm{v})$ can be pre-computed. $H\bm{q}$ only needs to be computed online via Fast Johnson–Lindenstrauss Transform ACHash ; falconn once, with cost $O(d\log d)$, and $\left\langle H\bm{q},Z_{S}(H\bm{v})\right\rangle$ can be computed online with cost $O(L)$ for every $\bm{v}$, where $L\ll d$ denotes the number of partitioned subspaces and will be explained in Sec. 3.2.
(2) (Exploitation of reference angle) In the design of existing projection techniques such as CEOs ceos , Falconn falconn , Falconn++ Falconn++ , etc., only the reference vector $Z_{S}{(\cdot))}$ is utilized. In contrast, our kernel functions defined in eq. (2) and eq. (3) incorporate not only the reference vector $Z_{S}{(\cdot))}$ but also the reference angle information $A_{S}(\cdot)$ (although the reference angle is not explicitly shown in eq. (2), its influence will become clear in Lemma 3.2). In fact, the reference angle plays a central role, as it is the key factor controlling the precision of angle estimation (see Lemma 3.2).
(3) (Generalizations of CEOs and PEOs) Beyond incorporating the reference angle, these two kernel functions can also be regarded as generalizations of CEOs and PEOs respectively in a certain sense. Specifically, if $S$ is taken as a point set generated via a Gaussian distribution and $Z_{S}(\bm{v})$ is replaced by the reference vector having the maximum inner product with the query, then $K^{1}_{S}(\bm{q},\bm{v})$ equals the indicator $\bm{v}^{\top}\bm{u_{\text{max}}}$ used in CEOs. Similarly, if we remove the term $A_{S}(H\bm{v})$ and take $S$ to be the same space-partitioned structure as that of PEOs, $K^{2}_{S}(\bm{q},\bm{v})$ is similar to the indicator of PEOs.
(4) (Configuration of projection vectors) Although we do not currently require any specific properties of the configuration of $S$, it is clear that the shape of $S$ impacts both $K^{1}_{S}(\bm{q},\bm{v})$ and $K^{2}_{S}(\bm{q},\bm{v})$. We will discuss the structure of $S$ in detail in Sec. 3.2.
Next, we provide a definition that will be used to establish the property of $K^{2}_{S}$.
Definition 3.1

### v2

@@ Abstract
In this paper, we study the angle testing problem in the context of similarity search in high-dimensional Euclidean spaces and propose two projection-based probabilistic kernel functions, one designed for angle comparison and the other for angle thresholding. Unlike existing approaches that rely on random projection vectors drawn from Gaussian distributions, our approach leverages reference angles and employs a deterministic structure for the projection vectors. Notably, our kernel functions do not require asymptotic assumptions, such as the number of projection vectors tending to infinity, and can be both theoretically and experimentally shown to outperform Gaussian-distribution-based kernel functions. We apply the proposed kernel function to Approximate Nearest Neighbor Search (ANNS) and demonstrate that our approach achieves a 2.5X $\sim$ 3X higher query-per-second (QPS) throughput compared to the widely-used graph-based search algorithm HNSW.
@@ 1 Introduction
Vector-based similarity search is a core problem with broad applications in machine learning, data mining, and information retrieval. It involves retrieving data points in a high-dimensional space that are most similar to a given query vector based on a specific similarity measure. This task is central to many downstream applications, including nearest neighbor classification, recommendation systems, clustering, anomaly detection, and retrieval-augmented generation (RAG). However, the high dimensionality of modern datasets makes efficient similarity search particularly challenging, highlighting the need for fast and scalable vector computation techniques.
@@
Among the various similarity measures for high-dimensional vectors, the $\ell_{2}$ norm, cosine distance, and inner product are the most commonly used in practice. As discussed in Yan et al. (2018); Dai et al. (2020); Lu et al. (2024), it is often possible to pre-compute and store the norms of vectors in advance, allowing these measures to be reduced to the computation of the cosine of the angle between two normalized vectors, thereby highlighting the central role of angle computation. On the other hand, in many real-world scenarios, we are not concerned with the exact values of the angles but rather with the outcome—which one is greater—of an angle comparison, which is referred to as the angle testing: Given a query vector $\bm{q}$ and data vectors $\bm{v_{1}}$, $\bm{v_{2}}$, $\bm{v}$ on sphere $\mathbb{S}^{d-1}$, typical operations include comparing $\left\langle\bm{q},\bm{v_{1}}\right\rangle$ and $\left\langle\bm{q},\bm{v_{2}}\right\rangle$, or determining whether $\left\langle\bm{q},\bm{v}\right\rangle$ exceeds a certain threshold. These operations, however, require computing exact cosines of angles, which have a cost of $O(d)$ per comparison and become expensive in high dimensions. To address this, we aim to design a computation-efficient probabilistic kernel function $K$ that can approximate these comparisons with reduced cost and high success probability. More precisely, we focus on the following two problems:
Problem 1.1.
@@
(Probabilistic kernel function for comparison) Design a probabilistic kernel function $K$: $\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}\rightarrow RV$ with computational cost $o(d)$, where $RV$ denotes the set of random variables, such that, for any data vectors $v_{1},v_{2}\in\mathbb{S}^{d-1}$ and query $q\in\mathbb{S}^{d-1}$ satisfying $\left\langle q,v_{1}\right\rangle>\left\langle q,v_{2}\right\rangle$, we have $\Pr[K(q,v_{1})>K(q,v_{2})]>1-\epsilon$, where $\epsilon\leq 0.5$.
Problem 1.2.
(Probabilistic kernel function for thresholding) Given a fixed angle threshold $\theta\in(0,\pi)$, design a probabilistic kernel function $K$: $\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}\rightarrow RV$ with computational cost $o(d)$ such that for any $q,v_{1},v_{2}\in\mathbb{S}^{d-1}$ with angles $\phi_{1}<\theta$ between $q$ and $v_{1}$, and $\phi_{2}>\theta$ between $q$ and $v_{2}$, we have $\Pr[K(q,v_{1})>\cos\theta]\geq 1-\epsilon_{1}$, and $\Pr[K(q,v_{2})>\cos\theta]<\epsilon_{2}$, where $\epsilon_{1},\epsilon_{2}\leq 0.5$.
@@
One of the main goals of this paper is to design appropriate $K$’s to solve the above two problems. Before proceeding, we note that much work has focused on the estimation of $\left\langle\bm{q},\bm{v}\right\rangle$, such as Johnson-Lindenstrauss (JL) bound (Johnson et al., 1984) and quantization-based techniques (Jégou et al., 2011; Ge et al., 2014; Martinez et al., 2018; Gao & Long, 2024; Guo et al., 2020). In contrast, the defined problems focus on comparing $\left\langle\bm{q},\bm{v}\right\rangle$ with another inner product or a threshold, necessitating a distinct theoretical analysis centered on probabilistic decision-making. Beyond the theoretical perspective, these two problems also give rise to a range of practical applications. The goal of Problem 1.1 aligns with that of the random-projection-based technique CEOs (Pham, 2021) under cosine distance (we postpone the general case of inner product to Sec. 6), allowing the designed kernel function to be applied to tasks where CEOs is effective, such as Maximum Inner Product Search (MIPS) (Pham, 2021), filtering of NN candidates (Pham & Liu, 2022), DBSCAN (Xu & Pham, 2024), and more. Notably, the goal of Problem 1.2 is similar to that of the graph-based ANNS approach PEOs (Lu et al., 2024), making the corresponding kernel function well suited for probabilistic routing tests in similarity graphs, which have demonstrated significant performance improvements over baseline graphs such as HNSW (Malkov & Yashunin, 2020). In Sec. 6, we will further elaborate on the applications of these two probabilistic kernel functions.
@@
Despite addressing different tasks, all the techniques (Pham, 2021; Pham & Liu, 2022; Xu & Pham, 2024; Lu et al., 2024) mentioned above use Gaussian distribution to generate projection vectors and are built upon a common statistical result, as follows.
Lemma 1.3.
@@
(Theorem 1 in Pham (2021)) Given two vectors $\bm{v}$, $\bm{q}$ on $\mathbb{S}^{d-1}$, and $m$ random vectors $\{\bm{u}_{i}\}^{m}_{i=1}\sim\mathcal{N}(0,I^{d})$, let $\bm{u}_{\mathrm{max}}={\mathop{\rm argmax}_{\bm{u_{i}}}}|\bm{q}^{\top}\bm{u_{i}}|$. As $m$ goes infinity, we have:
@@
However, Lemma 1.3 has a significant theoretical limitation: Relationship (1) relies on the assumption that the number of projection vectors $m$ tends to infinity. Since the evaluation time of projection vectors depends on $m$, $m$ cannot be very large in practice. Moreover, since Pham & Liu (2022); Xu & Pham (2024); Lu et al. (2024) all used Lemma 1.3 to derive their theoretical results, these results are also affected by this limitation, and the impact of $m$ becomes even harder to predict in these applications.
The starting point of our research is to overcome this limitation, and we make the following two key observations. (1) The Gaussian distribution used in Lemma 1.3 is not essential. Instead, the only factor determining the estimation accuracy of $\bm{q}^{\top}\bm{v}$ is the reference angle, that is, the angle between $\bm{q}$ and $\bm{u_{\text{max}}}$. (2) By introducing a random rotation matrix, the reference angle becomes dependent on the structure of the projection vectors and is predictable.
Based on these two observations, we design new probabilistic kernel functions to solve Problems 1.1 and 1.2. The contributions of this paper are summarized as follows.
(1) The proposed kernel functions $K^{1}_{S}$ and $K^{2}_{S}$ (Eq. (2) and Eq. (3)) rely on a reference-angle-based probabilistic relationship between angles in high-dimensional spaces and projected values. Compared with Eq. (1), the new relationship (Relationship (4)) is deterministic without dependence on asymptotic condition. By theoretical analysis, we show that the proposed kernel functions are effective solvers for the Problems 1.1 and 1.2 (see Lemmas 4.2,  4.3 and 5.1).
(2) By Lemmas 4.2,  4.3, we find that, the smaller the reference angle is, the more accurate the kernel functions are. To minimize the reference angle, we study the structure of the configuration of projection vectors (Sec. 5). We propose two structures (Alg. 1 and Alg. 2) that perform better than purely random projection (Alg. 3 in Appendix). We establish the relationship between the reference angle and the proposed structures (Lemma B.1 and Fig. 4 in Appendix).
(3) Based on $K^{1}_{S}$, we propose a random-projection technique KS1 which can be used for CEOs-based tasks (Pham, 2021; Pham & Liu, 2022; Xu & Pham, 2024). Based on $K^{2}_{S}$, we introduce a new routing test called the KS2 test which can be used to accelerate the graph-based Approximate Nearest Neighbor Search (ANNS) (Lu et al., 2024) (Sec. 6).
(4) We experimentally show that KS1 provides a slight accuracy improvement (up to 0.8%) over CEOs. For ANNS, we show that HNSW+KS2 improves the query-per-second (QPS) of the state-of-the-art approach HNSW+PEOs (Lu et al., 2024) by 10% – 30%, along with a 5% reduction in index size.
@@ 2 Related Work
Due to space limitations, we focus on random projection techniques that are closely related to this work. An illustration comparing the proposed random projection technique with others can be found in Sec. C.2 and Fig. 3 in Appendix. Since the proposed kernel function is also used in similarity graphs for ANNS, a comprehensive discussion of ANNS solutions is provided in Appendix C.3.
In high-dimensional Euclidean spaces, the estimation of angles via random-projection techniques, especially Locality Sensitive Hashing (LSH) (Indyk & Motwani, 1998; Andoni & Indyk, ; Andoni & Indyk, 2008), has a relatively long history. A classical LSH technique is SimHash (Charikar, 2002), whose basic idea is to generate multiple hyperplanes and partition the original space into many cells such that two vectors falling into the same cell are likely to have a small angle between them. Andoni et al. (2015) proposed a different LSH method called Falconn for angular distance, whose basic idea is to find the closest or furthest projection vector to the data vector and record this projection vector as a hash value, leading to better search performance than SimHash. Later, Pham (2021) employed Concomitants of Extreme Order Statistics (CEOs) to identify the projection with the largest or smallest inner product with the data vector, as shown in Lemma 1.3, and recorded the corresponding maximum or minimum projected value to obtain a more accurate estimation than using a hash value alone (Pham & Liu, 2022).
Due to its ease of implementation, CEOs has been employed in several similarity search tasks (Pham, 2021; Andoni et al., 2015; Xu & Pham, 2024), as mentioned in Sec. 1. Additionally, CEOs has been used to accelerate similarity graphs, which are among the leading structures for Approximate Nearest Neighbor Search (ANNS). By swapping the roles of query and data vectors in CEOs, Lu et al. (2024) introduced a space-partitioning technique and proposed the PEOs test, which can be used to compare the objective angle with a fixed threshold under probabilistic guarantees. This test was incorporated into the routing mechanisms of similarity graphs and achieved significant search performance improvements over original graph structures like HNSW (Malkov & Yashunin, 2020) and NSSG (Fu et al., 2022).
@@ 3 Two Probabilistic Kernel Functions
In Sec. 3, we aim to propose probabilistic kernel functions for Problems 1.1 and 1.2. First, we introduce some notation. Frequently used symbols in this paper are listed in Table 2 in Appendix. Let $\mathbb{R}^{d}$ be the ambient vector space. Define $H\in SO(d)\in\mathbb{R}^{d\times d}$ as a random rotation matrix [^1] Note that the definition here differs from that in Andoni et al. (2015), where the so-called random rotation matrix is actually a matrix with i.i.d. Gaussian entries) and let $S=[\bm{u_{1}},\bm{u_{2}},\ldots,\bm{u_{m}}]\in\mathbb{R}^{d\times m}$ be an arbitrary fixed set of $m$ points on the unit sphere $\mathbb{S}^{d-1}$. For any vector $\bm{v}\in\mathbb{S}^{d-1}$, define the reference vector of $\bm{v}$ with respect to $S$ as $Z_{S}(\bm{v})=\mathop{\rm argmax}_{\bm{u}\in S}\left\langle\bm{u},\bm{v}\right\rangle$. Let $A_{S}(\bm{v})$ denote the cosine of the reference angle with respect to $\bm{v}$, that is, $A_{S}(\bm{v})=\left\langle\bm{v},Z_{S}(\bm{v})\right\rangle$ . Next, we introduce two probabilistic kernel functions $K^{1}_{S}(\cdot,\cdot)$ and $K^{2}_{S}(\cdot,\cdot)$ as follows, where $K^{1}_{S}(\cdot,\cdot)$ corresponds to Problem 1.1 and $K^{2}_{S}(\cdot,\cdot)$ corresponds to Problem 1.2.
@@
Remarks. (1) (Exploitation of reference angle) In the design of existing projection techniques such as CEOs (Pham, 2021), Falconn (Andoni et al., 2015), Falconn++ (Pham & Liu, 2022), etc., only the reference vector $Z_{S}{(\cdot))}$ is utilized. In contrast, our kernel functions defined in Eq. (2) and Eq. (3) incorporate not only the reference vector $Z_{S}{(\cdot))}$ but also the reference angle information $A_{S}(\cdot)$ (although the reference angle is not explicitly shown in Eq. (2), its influence will become clear in Lemma 4.2). In fact, the reference angle plays a central role, as it is the key factor controlling the precision of angle estimation (see Lemma 4.2).
(2) (Generalizations of existing works) These two kernel functions can also be regarded as generalizations of CEOs and PEOs respectively in a certain sense. Specifically, if $S$ is taken as a point set generated via a Gaussian distribution and $Z_{S}(\bm{v})$ is replaced by the reference vector having the maximum inner product with the query, then $K^{1}_{S}(\bm{q},\bm{v})$ equals the indicator $\bm{v}^{\top}\bm{u_{\text{max}}}$ used in CEOs. Similarly, if we remove the term $A_{S}(H\bm{v})$ and take $S$ to be the same space-partitioned structure as that of PEOs, $K^{2}_{S}(\bm{q},\bm{v})$ is similar to the indicator of PEOs. In addition, Eq. (3) reduces to the estimator in Gao & Long (2024) when $S$ is taken to be the hypercube with scalar quantization.
(3) (Configuration of projection vectors) Although we do not currently require any specific properties of the configuration of $S$, it is clear that the configuration of $S$ impacts both $K^{1}_{S}(\bm{q},\bm{v})$ and $K^{2}_{S}(\bm{q},\bm{v})$. We will discuss the structure of $S$ in detail in Sec. 5. Notably, we will see that neither the hypercube adopted in Gao & Long (2024) nor the Gaussian distribution adopted in Pham (2021); Lu et al. (2024) provides the optimal configuration for projection vectors.
4 Analysis of Probability Guarantees
In Sec. 4, we show that the proposed probabilistic kernel functions $K^{1}_{S}$ and $K^{2}_{S}$ satisfy the probability guarantees of Problem 1.1 and Problem 1.2, respectively. Before proceeding, we first provide a definition that will be used to establish a property of $K^{2}_{S}$.
Definition 4.1.